Circular No. 7535

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000ez IN NGC 3995 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that spectra of SN 2000ez (cf. IAUC 7533), obtained by J. Huchra on Dec. 3.49 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) and by N. Caldwell on Dec. 3.53 with the MMT 6.5-m telescope (+ Blue Channel spectrograph), show it to be a type-II supernova several days after maximum light. The spectra consist of a continuum with superposed P-Cyg lines of H, He, and Ca II. Adopting the NASA/IPAC Extragalactic Database recession velocity of 3254 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 4000 km/s.

SUPERNOVA 2000fa IN UGC 3770 Matheson et al. also report that spectra of SN 2000fa, obtained as

above by Huchra on Dec. 3.46 UT and by Caldwell on Dec. 3.43, show it to be a type-Ia supernova well before maximum light. The spectra consist of a very blue continuum with well-defined Si II and Ca II features. Adopting the NASA/IPAC Extragalactic Database recession velocity of 6378 km/s for the host galaxy, the supernova expansion velocity is 16 000 km/s for Si II (rest 635.5 nm) and 27 000 km/s for Ca II (rest 395.1 nm). These expansion velocities imply an age of ∼ 2 weeks before maximum (following Jha et al. 1999, Ap.J. Supp. 125, 73).

Corrigenda. Regarding their discovery report on IAUC 7533, W. Li revises the position of SN 2000fa to be α = 7^{h}15^{m}29^{s}.88, δ = +23^{o}25′42′′.4 (equinox 2000.0), which is 6′′.6 east and 4′′.2 north of the nucleus of UGC 3770. Magnitudes for SN 2000fa: Nov. 30.5 UT, 17.3; Dec. 1.2, 17.0.

COMET C/2000 W1 (UTSUNOMIYA-JONES) Total visual magnitude and coma-diameter estimates: Nov. 27.50 UT,

7.1, 5′ (M. Mattiazzo, Wallaroo, S. Australia, 7×50 binoculars); 28.82, 7.2, 8′.0 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars); 30.53, 6.8, 8′.5 (Pearce); Dec. 1.91, 6.5, 5′ (P. M. Raymundo, Salvador, Brazil, 11×80 binoculars); 2.94, 6.8, 12′ (J. G. de S. Aguiar, Sousas, Brazil, 11×80 binoculars); 3.96, 6.9, 8′ (A. Amorim, Florianopolis, Brazil, 20×50 binoculars).

2000 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green